# E-SQL: Direct Schema Linking via Question Enrichment in Text-to-SQL

### v1

@@ 4.3. Schema Filtering
Schema filtering is a well-established technique aimed at minimizing the model’s reliance on irrelevant database schema elements by removing such items, and has been applied in numerous previous works. To evaluate the impact of schema filtering on a basic pipeline, we conducted various experiments both with and without a schema filtering module, which includes a filtered schema correction step. The results of these experiments are shown in Table 3. As demonstrated in the Table 3, the effect of schema filtering varies depending on its placement within the pipeline. Nevertheless, regardless of its position, incorporating the schema filtering module consistently led to a performance decline across all difficulty levels. Detailed results of the question enrichment module are discussed in the next section.
The inclusion of schema filtering resulted in an overall drop of up to 8.21% in Execution Accuracy (EX) on the development set. Specifically, the performance decreased by 6.38%, 12.28%, and 8.96% for simple, moderate, and challenging questions, respectively as shown in Table 3. Although the E-SQL pipeline does not inherently include a schema filtering module, we integrated it into the pipeline for our ablation study, as presented in Table 4. This integration resulted in a 5.54 drop in EX and a 5.68 decrease in Soft F1 score on the development set. These findings align with previous research, which suggests that advanced LLMs can manage schema linking effectively without requiring explicit filtering. Consequently, schema filtering was excluded from the final E-SQL pipeline, as its negative impact outweighed the potential benefits.
@@ 4.4. Ablation Study
We conducted an ablation study to analyze the contributions of the Question Enrichment, Candidate Predicate Augmentation, and SQL Refinement modules. The results of this study are summarized in Table 4, showing the impact individual modules from the E-SQL pipeline.
@@ 4.4.1. Question Enrichment
The Question Enrichment module, which facilitates direct schema linking by injecting database items, SQL components, conditions, and SQL generation steps into the question, improved performance across all difficulty levels, as shown in Table 4. Its impact was particularly significant on challenging questions. The absence of the question enrichment technique, especially when combined with the removal of the Candidate Predicate Augmentation module, led to a further decrease in performance. These results demonstrate that direct schema linking, achieved through question reformulation, effectively bridges the gap between the natural language query and the database schema, resulting in more accurate SQL generation.
@@ 4.4.2. Possible Predicate Augmentation
The Candidate Predicate Augmentation (CPG) module enhances the pipeline by augmenting potential predicates extracted from the database with the help of the LIKE operator and the candidate SQL query. As shown in the Table 4, removing the CPG module led to a nearly 2% drop in overall model performance. However, its removal slightly improved performance on challenging questions, suggesting that the CPG module may introduce unnecessary complexity in some cases.
@@ 4.4.3. SQL Refinement
The SQL Refinement (SR) module plays a crucial role in correcting minor errors in the generated SQL queries. Without SR, we observed a 3.57 drop in EX and a 3.73 decrease in Soft F1 across the development dataset. This demonstrates that the SQL refinement step significantly boosts the final query accuracy by detecting and correcting SQL execution errors.
@@
Table 5. Analysis of SQL Refinement (SR) Module on the Development Set
@@ 5. Discussion and Limitations
On the other hand, the schema filtering technique, though beneficial in prior research, led to a consistent performance decline even used with our novel question enrichment technique. The removal of this module resulted in performance improvements across all difficulty levels, reinforcing the conclusion that modern LLMs can grasp database schema without explicit filtering mechanisms.
@@
Despite the advancements, there are some limitations to our approach. Due to hardware and cost constraints, almost all experiments were conducted using GPT-4o-mini, leaving the impact of question enrichment and candidate predicate augmentation on open-source models unexplored. Investigating these techniques on open-source models remains an avenue for future work. Given the promising results with schema filtering in open-source LLMs, we anticipate that these techniques may further enhance performance in such settings.
@@ 6. Conclusions
In this study, we introduced E-SQL, a novel pipeline designed to address key challenges in Text-to-SQL translation by leveraging direct schema linking via question enrichment. Our experiments demonstrated that the question enrichment module, which augments natural language queries with relevant database elements and logical steps, significantly enhances query accuracy, particularly for complex queries. Furthermore, while prior works have highlighted the benefits of schema filtering, our findings reveal that its inclusion leads to performance degradation when using latest LLMs, supporting the notion that explicit schema filtering is redundant in modern architectures.
By excluding schema filtering and focusing on question enrichment and SQL refinement, E-SQL achieved competitive results on the BIRD benchmark. Specifically, E-SQL combined with GPT-4o achieved 65.58% and 66.29% execution accuracy on the development and test sets, respectively. These results underscore E-SQL’s effectiveness in handling complex queries and present it as a promising approach for future Text-to-SQL tasks.
Further exploration of open-source LLMs and iterative or multiple refinements are left for future work.
@@ Appendix B Manually Enriched Questions and Enrichment Reasoning
This section provides examples of manually enriched questions along with the rationale for their enrichment. Each manually enriched question and its corresponding reasoning are open-sourced and available at ¡githubLink¿.

### v2

@@ 4.3. Schema Filtering
Schema filtering is a technique aimed at minimizing the model’s reliance on irrelevant database schema elements by removing such items, and has been applied in numerous previous works. To evaluate the impact of schema filtering on a basic pipeline utilizing most advanced large language models, we conducted various experiments both with and without a schema filtering module, which includes a filtered schema correction step. The results of these experiments are shown in Table 4. As demonstrated in the Table 4, the effect of schema filtering varies depending on its placement within the pipeline. Nevertheless, regardless of its position, incorporating the schema filtering module in a basic pipeline consistently led to a performance decline across all difficulty levels. Detailed results of the question enrichment module are discussed in the next section.
The inclusion of schema filtering resulted in an overall drop of up to 8.21% in Execution Accuracy (EX) on the development set. Specifically, the performance decreased by 6.38%, 12.28%, and 8.96% for simple, moderate, and challenging questions, respectively as shown in Table 4. Although the E-SQL pipeline does not inherently include a schema filtering module, we integrated it into the pipeline for our ablation study, as presented in Table 5. This integration resulted in a 5.54 drop in EX and a 5.68 decrease in Soft F1 score on the development set. These findings align with previous research  (Maamari et al., 2024), which suggests that advanced LLMs can manage schema linking effectively without requiring explicit filtering. Consequently, schema filtering was excluded from the final E-SQL pipeline, as its negative impact outweighed the potential benefits.
@@ 4.4. Ablation Study
We conducted an ablation study to analyze the contributions of the Question Enrichment, Candidate Predicate Augmentation, and SQL Refinement modules. The results of this study are summarized in Table 5, showing the impact of individual modules from the E-SQL pipeline.
Table 6. Effect of QE and CPG modules in E-SQL pipeline with small open-source LLMs.
Pipeline Model Spider Test EX Bird Dev EX
E-SQL Qwen 2.5 Coder 7B 58.64 53.52
w/o QE Qwen 2.5 Coder 7B 55.84 $(\downarrow 2.80)$ 50.78 $(\downarrow 2.74)$
w/o CPG Qwen 2.5 Coder 7B 57.10 $(\downarrow 1.54)$ 50.72 $(\downarrow 2.80)$
w/o SR Qwen 2.5 Coder 7B 57.14 $(\downarrow 1.50)$ 48.24 $(\downarrow 5.28)$
@@ 4.4.1. Question Enrichment
The Question Enrichment module, which facilitates direct schema linking by injecting database items, SQL components, conditions, and SQL generation steps into the question, improved the performance as shown in Table 5 and Table 6 on both advance proprietary and small open-source LLMs. Its impact was particularly significant on challenging questions. The absence of the question enrichment technique, especially when combined with the removal of the Candidate Predicate Augmentation module (CPG), led to a further decrease in overall performance. These results demonstrate that direct schema linking, achieved through question reformulation, effectively bridges the gap between the natural language query and the database schema, resulting in more accurate SQL generation.
@@ 4.4.2. Possible Predicate Augmentation
The Candidate Predicate Augmentation (CPG) module enhances the pipeline by augmenting potential predicates extracted from the database with the help of the LIKE operator and the candidate SQL query. As shown in the Table 5 and Table 6, removing the CPG module resulted in a nearly 2% drop in overall model performance. However, its removal slightly improved the performance on challenging questions of BIRD dev set, suggesting that the CPG module may introduce unnecessary complexity in some cases. The slight negative effect of the CPG module on challenging questions is negligible, as it substantially enhances overall performance, especially when compared to the significant gains achieved through the Question Enrichment module.
@@ 4.4.3. SQL Refinement
The SQL Refinement (SR) module plays a crucial role in correcting minor errors in the generated SQL queries. Without SR, we observed a 3.57 drop in EX and a 3.73 decrease in Soft F1 across the BIRD development set. This demonstrates that the SQL refinement step significantly boosts the final query accuracy by detecting and correcting SQL execution errors.
@@
Table 7. Analysis of SQL Refinement (SR) Module on the BIRD Development Set
@@
4.5. Impact of Enriched Questions on Small Large Language Models
To evaluate the performance impact of database-integrated questions on small LLMs  (Yang et al., 2024; Hui et al., 2024; Guo et al., 2024), we conducted an experiment comparing their performance on default questions versus enriched questions. In this experiment, SQL queries for a given question were generated using a single-prompt approach. The prompt template used in this step is similar to that of the Candidate SQL Generation (CSG) module in the E-SQL pipeline, excluding data augmentation components such as database descriptions, value samples, and few-shot examples. This approach allows us to isolate the effect of enriched questions, as the single prompt relies solely on instructions and questions without additional context. Enriched questions were extracted from the outputs of Question Enrichment Module (QE) of the E-SQL pipeline utilized with GPT-4o to ensure that they incorporated database items and SQL construction plans.
The results, presented in  Table 8, indicate that even small LLMs can achieve competitive performance without task-specific fine-tuning when provided with high-quality, database-integrated natural language queries. These findings underscore the critical role of database-integrated enriched questions, which include logical SQL construction steps.
Table 8. Effect of Enriched Questions on the performance of small open-source LLMs without fine-tuning on BIRD development set. Enriched questions, generated using GPT-4o, were utilized to evaluate the impact of high-quality question enrichment on the performance of small open-source models. symbol denotes methods utilizing fine-tuned LLMs.
Model Level Dev EX
Default Enriched
DeepSeek Coder 1.3B Instruct Overall 20.92 50.84$(\uparrow 29.92)$
Simple 28.43 62.38$(\uparrow 33.95)$
Moderate 10.34 36.42$(\uparrow 26.08)$
Challenging 6.90 23.45$(\uparrow 16.55)$
Qwen2.5 Coder 1.5B Instruct Overall 11.21 36.90 $(\uparrow 25.69)$
Simple 14.27 44.10$(\uparrow 29.83)$
Moderate 6.68 28.23$(\uparrow 21.55)$
Challenging 6.20 18.62$(\uparrow 12.42)$
DeepSeek Coder 7B Instruct 1.5v Overall 37.02 56.45$(\uparrow 19.43)$
Simple 44.65 64.64$(\uparrow 19.99)$
Moderate 26.52 45.47$(\uparrow 18.95)$
Challenging 22.06 39.31$(\uparrow 17.25)$
Qwen2.5 Coder 7B Instruct Overall 31.25 40.22$(\uparrow 8.97)$
Simple 40.43 50.70$(\uparrow 10.27)$
Moderate 17.88 26.07$(\uparrow 8.19)$
Challenging 15.17 18.62$(\uparrow 3.45)$
ExSL + granite-20b-code Overall 51.69 -
DTS-SQL + DeepSeek 7B (Pourreza and Rafiei, 2024) Overall 55.8 -
SFT CodeS-7B  (Li et al., 2024c) Overall 57.17 -
SFT CodeS-15B  (Li et al., 2024c) Overall 58.47 -
4.6. Computational Cost Analysis
Understanding computational expenses is critical for assessing the practical scalability and applicability of the framework, especially given the reliance on large language models (LLMs) and their resource demands. Table 9 highlights the impact of question enrichment on token count2, while Table 10 provides details of average token usage for both prompt and completion stages of key pipeline components on the BIRD development set. The average number of tokens in prompts is inherently high due to well-defined instructions, data augmentations, including few-shot examples, database descriptions, and database values, as commonly employed in most Text-to-SQL approaches. Consequently, the increase in the question token count due to enrichment is relatively insignificant compared to the total number of prompt tokens. Despite the computational overhead introduced by question enrichment, which increases the average token count of natural language questions and their reasoning, E-SQL demonstrates computational superiority over methods  (Dong et al., 2023; Lee et al., 2024; Talaei et al., 2024; Gao et al., 2024; Maamari et al., 2024) that generate multiple SQL queries for a single user question. These methods incur significant computational costs due to the repeated SQL generation, subsequent correction of these SQL queries, and the selection process. Executing each E-SQL module only once minimizes the computational costs associated with repeated calls, ensuring greater efficiency. This balance between cost and performance underscores the scalability and efficiency of our pipeline for large-scale deployment.
22footnotetext: Token counts were computed using the tiktoken Python package, developed by OpenAI, which provides a programmatic interface for tokenizing text with OpenAI model-specific tokenizers. The package is available at [https://github.com/openai/tiktoken].
Table 9. Comparison of Default and Enriched Natural Language Questions in Bird Development Set
Text Avg. Tokens
Question 18.36
Enriched Question 81.51
Enrichment Reasoning 191.34
Fully Enriched Question 291.21
Table 10. Analysis of Computational Costs by Module on the Bird Development Set
Module Avg. Prompt Avg. Completion
Token Count Token Count
CSG 12612 199
QE 16550 292
SR 7403 267
In our pipeline, the initially generated candidate SQL query is executed to identify execution errors and enhance the large language model’s error awareness. While this step contributes to the overall pipeline latency, with an average execution time of 49.936 milliseconds per query, it plays a crucial role in ensuring accurate SQL refinement by providing valuable feedback on execution errors. It is important to note that the overall pipeline response time varies based on several factors, including the complexity of the natural language question, the length of the prompt, and the API response time of proprietary LLMs, which is influenced by server load and volume. These factors collectively contribute to the latency of the system.
@@ 5. Discussion and Limitations
The prompt design plays a critical role in influencing model performance. The prompt templates utilized for each module of the E-SQL pipeline are publicly available in our GitHub repository. This study primarily emphasizes schema linking through question enrichment and data augmentation, deliberately leaving the exploration of alternative prompt templates beyond its scope.
Despite the advancements, there are some limitations to our approach. Due to hardware and cost constraints, almost all experiments were conducted using GPT-4o-mini and small open-source LLMs without fine-tuning. Among the small open-source LLMs, the whole E-SQL pipeline was executed only with Qwen2.5 Coder 7B Instruct since the context length of the other small LLMs is not sufficient to run and observe the effect of E-SQL pipeline. Developing more efficient schema linking techniques that operate effectively with small LLMs and limited context lengths represents a promising direction for future work.
@@ 6. Conclusions
In this study, we introduced E-SQL, a novel pipeline designed to address key challenges in Text-to-SQL translation by leveraging direct schema linking via question enrichment and incorporating candidate predicates. Our experiments demonstrated that the question enrichment module, which integrates natural language queries with relevant database elements and logical steps, significantly enhances query accuracy, particularly for complex queries. Additionally, the proposed candidate predicate augmentation technique further improves the performance of the pipeline. Moreover, our additional experiments reveals the importance and positive impact of enriched questions on the performance of small open-source LLMs with limited context lengths.
While some prior works have highlighted the utility of schema filtering, our findings reveal that incorporating schema filtering into a text-to-SQL translation pipeline that leverages advanced LLMs results in performance degradation. This supports the notion that explicit schema filtering can be redundant in modern architectures that utilizes the latest LLMs.
By focusing on question enrichment, data augmentation and SQL refinement, E-SQL achieved competitive results on the BIRD benchmark. Specifically, E-SQL combined with GPT-4o achieved 65.58% and 66.29% execution accuracy on the development and test sets, respectively. These results underscore E-SQL’s effectiveness in handling complex queries and present it as a promising approach for future Text-to-SQL tasks.
Despite a minor computational overhead due to increased token counts in question enrichment, its impact is negligible compared to the overall token usage and is outweighed by the significant performance gains. Additionally, E-SQL ensures cost-efficiency by executing each module only once, avoiding the excessive resource demands of repeated query generation and correction. This balance highlights E-SQL’s scalability and suitability for resource-constrained deployments.
Further exploration of fine tuning, iterative or multiple question refinements and schema linking techniques optimized for small LLMs with limited context lengths is left for future work.
@@ Appendix B Manually Enriched Questions and Enrichment Reasoning
This section provides examples of manually enriched questions along with the rationale for their enrichment. Each manually enriched question and its corresponding reasoning are open-sourced and available at Github repository.